### v1

@@ 4.3 Necessary and Sufficient Conditions for These Codes to be Dual-Containing
Furthermore, since 00 is not a root of $x^{m}-1$, and hence not a root of $g_{1}^{*}(x)$, we conclude that $g_{1}^{*}(x)\mid\overline{g_{1}^{\perp}}(x)(1+v_{2}(x)\overline{v_{2}}(x))$ is equivalent to $g_{1}(x)\mid g_{1}^{\perp}(x)(1+\overline{v_{2}}(x)v_{2}(x))$. ∎
Lemma 4.
Lv2020-2 Let $f(x)$, $g(x)$, and $k(x)$ be polynomials over $\mathcal{R}$ (Here the base field is $\mathbb{F}_{q^{2}}$). Then, the following property holds for the Hermitian inner product:
Proposition 10.
Let $\mathcal{C}_{q^{2}}$ be the QC code in Definition 4. Then its Hermitian dual code $\mathcal{C}_{q^{2}}^{\perp_{H}}$ is the QC code generated by $([g_{1}^{\perp_{H}}(x)],[-\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)])$ and $([-\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)],[g_{2}^{\perp_{H}}(x)])$.
Proof.
Similar to the proof of Proposition 8, let $\mathcal{D}$ denote the QC code generated by $([g_{1}^{\perp_{H}}(x)],[-\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)])$ and $([-\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)],[g_{2}^{\perp_{H}}(x)])$. Then $\gcd(\overline{v}_{1}^{[q]}(x)\overline{v}_{2}^{[q]}(x)-1,x^{m}-1)=1$ follows from equation $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$, which immediately implies that $\dim(\mathcal{D})=\deg(g_{1})+\deg(g_{2}(x))$. Let $\mathbf{c}$ be an arbitrary codeword in $\mathcal{C}_{q^{2}}$ and
$\mathbf{c}=([a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)]),$
where $a_{1}(x),a_{2}(x)$ in $\mathbb{F}_{q^{2}}[x]/(x^{m}-1)$. Let $\mathbf{d}$ be an arbitrary codeword in $\mathcal{D}$ and write it as
$\mathbf{d}=([b_{1}(x)g_{1}^{\perp_{H}}(x)-b_{2}(x)\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)],[-b_{1}(x)\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)+b_{2}(x)g_{2}^{\perp_{H}}(x)]),$
where $b_{1}(x)$, $b_{2}(x)$ in $\mathbb{F}_{q^{2}}[x]/(x^{m}-1)$. The Hermitian inner product between $\mathbf{c}$ and $\mathbf{d}$ is given by
$\langle\mathbf{c},$ $\mathbf{d}\rangle_{H}=\langle[a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[b_{1}(x)g_{1}^{\perp_{H}}(x)-b_{2}(x)\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)]\rangle_{H}$
$+\langle[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)],[-b_{1}(x)\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)+b_{2}(x)g_{2}^{\perp_{H}}(x)]\rangle_{H}$
$=\langle[a_{1}(x)g_{1}(x)],[-b_{2}(x)\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)]\rangle_{H}+\langle[a_{2}(x)v_{2}(x)g_{2}(x)],[b_{1}(x)g_{1}^{\perp_{H}}(x)]\rangle_{H}$
$+\langle[a_{1}(x)v_{1}(x)g_{1}(x)],[b_{2}(x)g_{2}^{\perp_{H}}(x)]\rangle_{H}-\langle[a_{2}(x)g_{2}(x)],[b_{1}(x)\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)]\rangle_{H}.$
By Lemma 4, we know that the above equation is equal to 0. Thus, we know that $\mathcal{D}\subseteq\mathcal{C}_{q^{2}}^{\perp_{H}}$. Then we have $\mathcal{D}=\mathcal{C}_{q^{2}}^{\perp_{H}}$ from $\dim(\mathcal{D})=\dim(\mathcal{C}_{q^{2}}^{\perp_{H}})$, which completes the proof. ∎
Theorem 11.
The QC code $\mathcal{C}_{q^{2}}$ in Definition 4 is Hermitian dual-containing if and only if all the following conditions hold simultaneously:
$(1)\enspace g_{1}(x)\mid g_{1}^{\perp_{H}}(x)(1+v_{2}(x)\overline{v}_{2}^{[q]}(x)),$
$(2)\enspace g_{1}(x)\mid g_{2}^{\perp_{H}}(x)(\overline{v}_{1}^{[q]}(x)+v_{2}(x)),$
$(3)\enspace g_{2}(x)\mid g_{2}^{\perp_{H}}(x)(1+v_{1}(x)\overline{v}_{1}^{[q]}(x)).$
Proof.
Note that $x^{m}-1=\left(g_{1}^{\perp_{H}}\right)^{[q]}\left(\frac{x^{m}-1}{g_{1}^{\perp_{H}}}\right)^{[q]}$ and
$\left(\frac{x^{m}-1}{g_{1}^{\perp_{H}}(x)}\right)^{[q]}=\left(\frac{x^{m}-1}{\left(g_{1}^{\perp}(x)\right)^{[q]}}\right)^{[q]}=\frac{x^{m}-1}{g_{1}^{\perp}(x)}=g_{1}^{*}(x).$
Similar to the proof of Theorem 9, we can conclude that this class of codes is Hermitian dual-containing if and only if $g_{1}^{*}(x)\mid\overline{g_{1}^{\perp_{H}}(x)}(1+\overline{v_{2}}(x)v_{2}^{[q]}(x))$, $g_{1}^{*}(x)\mid\overline{g_{2}^{\perp_{H}}(x)}(v_{1}^{[q]}(x)+\overline{v_{2}}(x))$, and $g_{1}^{*}(x)\mid\overline{g_{1}^{\perp_{H}}(x)}(1+\overline{v_{1}}(x)v_{1}^{[q]}(x))$ from Proposition 10 and Theorem 2. Again, by a similar argument as in Theorem 9, these conditions are essentially equivalent to the three conditions stated in the current theorem. ∎
Proposition 12.
Let $\mathcal{C}$ be the QC code in Definition 4. Then its Symplectic dual code $\mathcal{C}^{\perp_{S}}$ is the QC code generated by $([g_{2}^{\perp}(x)],[\overline{v}_{1}(x)g_{2}^{\perp}(x)])$ and $([\overline{v}_{2}(x)g_{1}^{\perp}(x)],[g_{1}^{\perp}(x)])$.
Proof.
Similar to the proof of Proposition 8, let $\mathcal{D}$ denote the QC code generated by $([g_{2}^{\perp}(x)],[\overline{v}_{1}(x)g_{2}^{\perp}(x)])$ and $([\overline{v}_{2}(x)g_{1}^{\perp}(x)],[g_{1}^{\perp}(x)])$. Then $\dim(\mathcal{D})=\deg(g_{1})+\deg(g_{2})$. Let $\mathbf{c}$ be an arbitrary codeword in $\mathcal{C}$ and
$\mathbf{c}=([a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)]),$
where $a_{1}(x)$,$a_{2}(x)$ in $\mathcal{R}$. Let $\mathbf{d}$ be an arbitrary codeword in $\mathcal{D}$ and write it as
$\mathbf{d}=([b_{1}(x)g_{2}^{\perp}(x)+b_{2}(x)\overline{v}_{2}(x)g_{1}^{\perp}(x)],[b_{1}(x)\overline{v}_{1}(x)g_{2}^{\perp}(x)+b_{2}(x)g_{1}^{\perp}(x)]),$
where $b_{1}(x)$, $b_{2}(x)$ in $\mathcal{R}$. The Symplectic inner product between $\mathbf{c}$ and $\mathbf{d}$ is given by
$\langle\mathbf{c},\mathbf{d}\rangle_{S}$ $=\langle{[a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[b_{1}(x)\overline{v}_{1}(x)g_{2}^{\perp}(x)+b_{2}(x)g_{1}^{\perp}(x)]}\rangle_{E}$
$\quad-\langle{[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)],[b_{1}(x)g_{2}^{\perp}(x)+b_{2}(x)\overline{v}_{2}(x)g_{1}^{\perp}(x)]}\rangle_{E}$
$=\langle{[a_{1}(x)g_{1}(x)],[b_{1}(x)\overline{v}_{1}(x)g_{2}^{\perp}(x)}]\rangle_{E}+\langle{[a_{2}(x)v_{2}(x)g_{2}(x)],[b_{2}(x)g_{1}^{\perp}(x)]}\rangle_{E}$
$\quad-\langle{[a_{1}(x)v_{1}(x)g_{1}(x)],[b_{1}(x)g_{2}^{\perp}(x)]}\rangle_{E}-\langle{[a_{2}(x)g_{2}(x)],[b_{2}(x)\overline{v}_{2}(x)g_{1}^{\perp}(x)]}\rangle_{E}.$
According to Lemma 3, it can be verified that the above equation is equal to 00. Therefore, $\mathcal{D}\subseteq\mathcal{C}^{\perp_{S}}$. Hence, we get $\mathcal{D}=\mathcal{C}^{\perp_{S}}$ from $\dim(\mathcal{D})=\dim(\mathcal{C}^{\perp_{S}})$. The proof is complete. ∎
Theorem 13.
The QC code defined in Definition 4 is symplectic dual-containing if and only if all the following conditions hold simultaneously.
$(1)\enspace g_{2}(x)\mid g_{2}^{\perp}(x)(\overline{v_{1}}(x)-v_{1}(x)),$
$(2)\enspace g_{2}(x)\mid g_{1}^{\perp}(x)(1-\overline{v_{1}}(x)v_{2}(x)),$
$(3)\enspace g_{1}(x)\mid g_{1}^{\perp}(x)(v_{2}(x)-\overline{v_{2}}(x)).$
Proof.
Similar to the proof of Theorem 9, we can conclude that this class of codes is symplectic dual-containing if and only if $g_{2}^{*}(x)\mid\overline{g_{2}^{\perp}}(x)(v_{1}(x)-\overline{v_{1}}(x))$, $g_{2}^{*}(x)\mid\overline{g_{1}^{\perp}}(x)(1-v_{1}(x)\overline{v_{2}}(x))$ and $g_{1}^{*}(x)\mid\overline{g_{1}^{\perp}}(x)(\overline{v_{2}}(x)-v_{2}(x))$ from Proposition 12 and Theorem 3. Again, by a similar argument as in Theorem 9, these conditions are essentially equivalent to the three conditions stated in the current theorem. ∎
@@ 4.4.1 QC Codes for $v_{2}(x)=0$ in Definition 4
By substituting $v_{2}(x)=0$ into Theorems 9, 11, and 13, we get the following three corollaries.
@@
The QC code $\mathcal{D}_{1}$ in Definition 5 is Euclidean dual-containing if and only if all the following conditions hold simultaneously.
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp}(x)\mid g_{1}^{\perp}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be Euclidean dual-containing given in Proposition 11 of Galindo2018 , is a special case of the necessary and sufficient conditions in Corollary 1. This can also be seen from the following example.
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp_{H}}(x)\mid g_{1}^{\perp_{H}}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be Hermitian dual-containing given in Proposition 15 of Galindo2018 , is a special case of the necessary and sufficient conditions in Corollary 2. This can also be seen from the following example.
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp}(x)\mid g_{1}^{\perp}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be symplectic dual-containing given in Theorem 5 (case (ii)) of Galindo2018 , is a special case of the necessary and sufficient conditions in Corollary 3. This can also be seen from the following example.
@@ 4.4.2 QC Codes for $v_{2}(x)=1$ in Definition 4
By substituting $v_{2}(x)=1$ into Theorems 9 and 13, we get the following corollaries.
@@
The conditions $g_{2}(x)\mid g_{1}^{\perp}(x)$, $g_{2}(x)\mid g_{2}^{\perp}(x)$, and $\overline{v_{1}}(x)=v_{1}(x)$, as sufficient conditions for $\mathcal{D}_{2}$ to be Euclidean dual-containing given in Lemma 10 (It is valid only for $q=2$) of duchao2023 , are a special case of the necessary and sufficient conditions in Corollary 4 (It is valid for any prime power $q$) because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@
The conditions $g_{1}(x)\mid g_{2}^{\perp}(x)$ and $\overline{v_{1}}(x)=v_{1}(x)$, as sufficient conditions for $\mathcal{D}_{2}$ to be symplectic dual-containing stated in Proposition 7 of Lv2020 , are a special case of the necessary and sufficient conditions in Corollary 5 because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@ 4.4.3 QC Codes for $v_{1}(x)=v_{2}(x)$ in Definition 4
For the definition of the QC code $\mathcal{D}_{3}$, Proposition 1 of Guan2022-2 requires that $v(x)$ (denoted as $t(x)$ there) is a polynomial over $\mathcal{R}_{2}=\mathbb{F}_{q^{2}}/(x^{m}-1)$ such that $\dim\mathcal{D}_{3}=2m-\deg(g_{1}(x))-\deg(g_{2}(x))$, but does not explicitly specify constraints on it. In our approach, we impose the condition $\gcd(v^{2}(x)-1,x^{m}-1)=1$, ensuring that $\dim\mathcal{D}_{3}=2m-\deg(g_{1}(x))-\deg(g_{2}(x))$.
By substituting $v_{1}(x)=v_{2}(x)=v(x)$ into theorems 11 and 13, we get the following corollaries.
@@
The conditions $g_{1}(x)\mid g_{1}^{\perp_{H}}(x)$, $g_{1}(x)\mid g_{2}^{\perp_{H}}(x)(\overline{v}^{[q]}(x)+v(x))$, and $g_{2}(x)\mid g_{2}^{\perp_{H}}(x)$, as sufficient conditions for $\mathcal{D}_{3}$ to be Hermitian dual-containing stated in Proposition 2 of Guan2022-2 , are special cases of the necessary and sufficient conditions in Corollary 6. This can also be seen from the following example.
@@
The conditions $g_{1}(x)\mid g_{2}^{\perp}(x)$, $\gcd(g_{1}(x),g_{2}(x))=1$, and $\overline{v}(x)=v(x)$, as sufficient conditions for $D_{3}$ to be symplectic dual-containing stated in Proposition 4 (It is valid only for $q=2$) of  Guan2022-1 , are special cases of the necessary and sufficient conditions in Corollary  7 (It is valid for any prime power $q$) because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@ 5 Construction of Quantum Stabilizer Codes
This section extends the results established in the previous section to construct quantum stabilizer codes from self-orthogonal QC codes. With the aid of the algebraic computation software Magma magma , we systematically compute the parameters of these codes, some of which match the best-known quantum error-correcting codes listed in Grassl’s code tables codetables . These results highlight the potential of QC codes in constructing good quantum codes.
To begin with, we provide the following lemmas.
@@
Ashikhmin2001 Given two classical linear codes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ of length $n$ over $\mathbb{F}_{q}$ with $\mathcal{C}_{2}\subset\mathcal{C}_{1}$, we can construct a stabilizer quantum code with parameters
$[[n,\dim\mathcal{C}_{1}-\dim\mathcal{C}_{2},\min\{w(\mathcal{C}_{1}\setminus\mathcal{C}_{2}),w(\mathcal{C}_{2}^{\perp}\setminus\mathcal{C}_{1}^{\perp})\}]]_{q}.$
@@
Ketkar2006 Let $\mathcal{C}$ be a Hermitian self-orthogonal linear code $[n,k]_{q^{2}}$ over $\mathbb{F}_{q^{2}}$, such that every nonzero vector in $\mathcal{C}^{\perp_{H}}\setminus\mathcal{C}$ has weight at least $d^{\prime}$. Then, $\mathcal{C}$ induces a quantum stabilizer code with parameters
$[[n,n-2k,d^{\prime}]]_{q}.$
@@
Calderbank1998 Let $\mathcal{C}$ be a symplectic self-orthogonal $[2n,k]$ linear code over $\mathbb{F}_{q}$. Then, a quantum stabilizer code with parameters
$[[n,n-k,d^{\prime}]]_{q}$
@@
According to Theorem 5 and Lemma 5, a QC code satisfying the Euclidean self-orthogonal condition can be employed to construct a corresponding quantum stabilizer code, as detailed below.
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a Euclidean self-orthogonal code with parameters $[2m,2m-\deg(g_{1})-\deg(g_{2})]$. Therefore, there exists a quantum stabilizer code with the following parameters:
$[[2m,2(\deg(g_{1})+\deg(g_{2})-m),w(\mathcal{C}^{\perp}\setminus\mathcal{C})]]_{q}.$
@@
Let $q=2$, $m=9$, $g_{1}(x)=x^{7}+x^{6}+x^{4}+x^{3}+x+1$, $g_{2}(x)=x^{8}+x^{7}+x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1$, $v_{1}(x)=x^{8}+x+1$, $v_{2}(x)=x^{8}+1$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{E}}$ are $[18,14,2]_{2}$. Applying Theorem 14, we construct a quantum error-correcting code with parameters $[[18,12,2]]_{2}$, which coincide with those of a best-known code listed in Grassl’s code table codetables .
@@
Suppose that $\mathcal{C}_{q^{2}}$ is a 2-generator QC code defined in Definition 4 and satisfies the following conditions:
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a Hermitian self-orthogonal code with parameters $[2m,2m-\deg(g_{1})-\deg(g_{2})]_{q^{2}}$. Therefore, there exists a quantum stabilizer code with the following parameters:
$[[2m,2(\deg(g_{1})+\deg(g_{2})-m),w(\mathcal{C}^{\perp_{H}}\setminus\mathcal{C})]]_{q}.$
@@
Let $q^{2}=4$, $m=3$, $\mathbb{F}_{4}=\mathbb{F}_{2}(w)$, $g_{1}(x)=x^{2}+wx+w+1$, $g_{2}(x)=x^{2}+(w+1)x+w$, $v_{1}(x)=x^{2}$, $v_{2}(x)=x^{2}+w+1$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{H}}$ are $[6,2,2]_{2}$. Applying Theorem 15, we construct a quantum error-correcting code with parameters $[[6,2,2]]_{2}$, which coincide with those of a best-known code listed in Grassl’s code table codetables .
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a symplectic self-orthogonal code with parameters $[2m,2m-\deg(g_{1})-\deg(g_{2})]$. Therefore, there exists a quantum stabilizer code with the following parameters:
$[[m,\deg(g_{1})+\deg(g_{2})-m,w_{s}(\mathcal{C}^{\perp_{S}}\setminus\mathcal{C})]]_{q}.$

### v2

@@ 4.1 Definition and analysis of the new QC codes
Furthermore, 0 is not a root of $x^{m}-1$, so it is also not a root of $g_{1}^{*}(x)$. Therefore, we conclude that $g_{1}^{*}(x)\mid\overline{g_{1}^{\perp}}(x)(1+v_{2}(x)\overline{v_{2}}(x))$ is equivalent to $g_{1}(x)\mid g_{1}^{\perp}(x)(1+\overline{v_{2}}(x)v_{2}(x))$. The result for the Euclidean inner product case follows. By similar arguments, the results for the remaining two inner products can also be obtained and are therefore omitted. ∎
@@ 4.2.1 QC codes for $v_{2}(x)=0$ in Definition 4
By substituting $v_{2}(x)=0$ in Theorem 5, we get the following three corollaries.
@@
The QC code $\mathcal{D}_{1}$ in Definition 5 is Euclidean dual-containing if and only if all the following conditions hold simultaneously:
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp}(x)\mid g_{1}^{\perp}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be Euclidean dual-containing given in Proposition 11 of [24], is a special case of the necessary and sufficient conditions in Corollary 1. This can also be seen from the following example.
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp_{H}}(x)\mid g_{1}^{\perp_{H}}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be Hermitian dual-containing given in Proposition 15 of [24], is a special case of the necessary and sufficient conditions in Corollary 2. This can also be seen from the following example.
@@
The condition $g_{1}(x)\mid g_{2}(x)\mid g_{2}^{\perp}(x)\mid g_{1}^{\perp}(x)$, as a sufficient condition for $\mathcal{D}_{1}$ to be symplectic dual-containing given in Theorem 5 (case (ii)) of [24], is a special case of the necessary and sufficient conditions in Corollary 3. This can also be seen from the following example.
@@ 4.2.2 QC codes for $v_{2}(x)=1$ in Definition 4
Substituting $v_{2}(x)=1$ into Cases A and C of Theorem 5 yields the following corollaries.
@@
The conditions $g_{2}(x)\mid g_{1}^{\perp}(x)$, $g_{2}(x)\mid g_{2}^{\perp}(x)$, and $\overline{v_{1}}(x)=v_{1}(x)$, as sufficient conditions for $\mathcal{D}_{2}$ to be Euclidean dual-containing given in Lemma 10 (It is valid only for $q=2$) of [34], are a special case of the necessary and sufficient conditions in Corollary 4 (It is valid for any prime power $q$) because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@
The conditions $g_{1}(x)\mid g_{2}^{\perp}(x)$ and $\overline{v_{1}}(x)=v_{1}(x)$, as sufficient conditions for $\mathcal{D}_{2}$ to be symplectic dual-containing stated in Proposition 7 of [25], are a special case of the necessary and sufficient conditions in Corollary 5 because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@ 4.2.3 QC codes for $v_{1}(x)=v_{2}(x)$ in Definition 4
For the definition of the QC code $\mathcal{D}_{3}$, Proposition 1 of [28] requires that $v(x)$ (denoted as $t(x)$ there) is a polynomial over $\mathcal{R}_{2}=\mathbb{F}_{q^{2}}/(x^{m}-1)$ such that $\dim\mathcal{D}_{3}=2m-\deg(g_{1}(x))-\deg(g_{2}(x))$, but does not explicitly specify constraints on it. In our approach, We have included a concrete formulation of the condition, i.e., $\gcd(v^{2}(x)-1,x^{m}-1)=1$, ensuring that $\dim\mathcal{D}_{3}=2m-\deg(g_{1}(x))-\deg(g_{2}(x))$.
The following corollaries follow from Cases B and C of Theorem 5 by setting $v_{1}(x)=v_{2}(x)=v(x)$.
@@
The conditions $g_{1}(x)\mid g_{1}^{\perp_{H}}(x)$, $g_{1}(x)\mid g_{2}^{\perp_{H}}(x)(\overline{v}^{[q]}(x)+v(x))$, and $g_{2}(x)\mid g_{2}^{\perp_{H}}(x)$, as sufficient conditions for $\mathcal{D}_{3}$ to be Hermitian dual-containing stated in Proposition 2 of [28], are special cases of the necessary and sufficient conditions in Corollary 6. This can also be seen from the following example.
@@
The conditions $g_{1}(x)\mid g_{2}^{\perp}(x)$, $\gcd(g_{1}(x),g_{2}(x))=1$, and $\overline{v}(x)=v(x)$, as sufficient conditions for $D_{3}$ to be symplectic dual-containing stated in Proposition 4 (It is valid only for $q=2$) of  [29], are special cases of the necessary and sufficient conditions in Corollary  7 (It is valid for any prime power $q$) because $g_{1}(x)\mid g_{2}^{\perp}(x)\iff g_{2}(x)\mid g_{1}^{\perp}(x)$. This can also be seen from the following example.
@@ 5 Construction of quantum stabilizer codes
This section extends the results established in the previous section to construct quantum stabilizer codes from self-orthogonal QC codes. With the aid of the algebraic computation software Magma  [37], we systematically compute the parameters of these codes, some of which match the best-known quantum error-correcting codes listed in the codetable of Grassl [36]. These results highlight the potential of QC codes in constructing good quantum codes. To begin with, we provide the following lemmas.
@@
[8] Given two classical linear codes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ of length $n$ over $\mathbb{F}_{q}$ with $\mathcal{C}_{2}\subset\mathcal{C}_{1}$, we can construct a stabilizer quantum code with parameters
$n,\dim\mathcal{C}_{1}-\dim\mathcal{C}_{2},\min\{w(\mathcal{C}_{1}\setminus\mathcal{C}_{2}),w(\mathcal{C}_{2}^{\perp}\setminus\mathcal{C}_{1}^{\perp})\}]]_{q}.$
@@
[9] Let $\mathcal{C}$ be a Hermitian self-orthogonal linear code $n,k]_{q^{2}}$ over $\mathbb{F}_{q^{2}}$, such that every nonzero vector in $\mathcal{C}^{\perp_{H}}\setminus\mathcal{C}$ has weight at least $d^{\prime}$. Then, $\mathcal{C}$ induces a quantum stabilizer code with parameters
$n,n-2k,d^{\prime}]]_{q}.$
@@
[7] Let $\mathcal{C}$ be a symplectic self-orthogonal $[2n,k]$ linear code over $\mathbb{F}_{q}$. Then, a quantum stabilizer code with parameters
$n,n-k,d^{\prime}]]_{q}$
@@
According to Theorem 3 and Lemma 5, a quasi-cyclic (QC) code that satisfies the Euclidean self-orthogonality condition can be used to construct a corresponding quantum stabilizer code, as detailed below.
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a Euclidean self-orthogonal code with parameters $2m,2m-\deg(g_{1})-\deg(g_{2})]$. Therefore, there exists a quantum stabilizer code with the following parameters:
$2m,2(\deg(g_{1})+\deg(g_{2})-m),w(\mathcal{C}^{\perp}\setminus\mathcal{C})]]_{q}.$
@@
Let $q=2$, $m=9$, $g_{1}(x)=x^{7}+x^{6}+x^{4}+x^{3}+x+1$, $g_{2}(x)=x^{8}+x^{7}+x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1$, $v_{1}(x)=x^{8}+x+1$, $v_{2}(x)=x^{8}+1$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{E}}$ are $18,14,2]_{2}$. Applying Theorem 6, we construct a quantum error-correcting code with parameters $18,12,2]]_{2}$, which coincide with those of a best-known code listed in the codetable of Grassl [36].
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a Hermitian self-orthogonal code with parameters $2m,2m-\deg(g_{1})-\deg(g_{2})]_{q^{2}}$. Therefore, there exists a quantum stabilizer code with the following parameters:
$2m,2(\deg(g_{1})+\deg(g_{2})-m),w(\mathcal{C}^{\perp_{H}}\setminus\mathcal{C})]]_{q}.$
@@
Let $q^{2}=4$, $m=3$, $\mathbb{F}_{4}=\mathbb{F}_{2}(w)$, $g_{1}(x)=x^{2}+wx+w+1$, $g_{2}(x)=x^{2}+(w+1)x+w$, $v_{1}(x)=x^{2}$, $v_{2}(x)=x^{2}+w+1$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{H}}$ are $6,2,2]_{2}$. Applying Theorem 7, we construct a quantum error-correcting code with parameters $6,2,2]]_{2}$, which coincide with those of a best-known code listed in the codetable of Grassl [36].
@@
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then, $\mathcal{C}$ is a symplectic self-orthogonal code with parameters $2m,2m-\deg(g_{1})-\deg(g_{2})]$. Therefore, there exists a quantum stabilizer code with the following parameters:
$m,\deg(g_{1})+\deg(g_{2})-m,w_{s}(\mathcal{C}^{\perp_{S}}\setminus\mathcal{C})]]_{q}.$